PROBLEM 16-1 Simple Liquidation
The Discount Partnership is being liquidated. The current balance sheet is shown here.
Discount Partnership
Balance Sheet
14-Jan-14
Assets
Cash $ 25,000
Other assets 120,000
Total assets $145,000
Liabilities and Partners’ Equity
Accounts payable $ 40,000
Dawson, capital 31,000
Feeney, capital 65,000
Hardin, capital 9,000
Total liabilities and partners’ equity $145,000
Dawson, Feeney, and Hardin share profits and losses in a 30:40:30 ratio.

Required:

A. Prepare a schedule of partnership liquidation for each of the following three independent cases.
(1) The noncash assets are sold for $60,000, and any partner with a deficit is unable to eliminate any of the
deficit.
(2) The noncash assets are sold for $60,000, and any partner with a deficit is able to invest cash equal to the
amount of the deficit.
(3) The noncash assets are sold for $50,000, and any partner with a deficit is able to invest up to $8,000 cash
in the partnership.
B. Prepare all necessary journal entries for case 2 above.
Answer #1
1st Case
The noncash assets are sold for $60,000, and any partner with a deficit is unable to eliminate any of the deficit.
cash available Dawson Feeney Hardin
$45,000 $13,000 $41,000 ($9,000)**
0 ($3,857) = $9,000*30/70 (5143) = $9,000*40/70 $9,000
Available balance =$45,000 $9,143 $35,857 0
cash distributed to surplus partner($45,000) ($9,143) ($35,857) 0
balance 0 0 0 0
2nd Case The noncash assets are sold for $60,000, and any partner with a deficit is able to invest cash equal to the amount of the deficit.
cash balance Dawson Feeney Hardin
$45,000 $13,000 $41,000 ($9,000)
Deficient partner will meet its deficiency =$9,000 0 0 $9,000
available balance =$54,000 $13,000 $41,000 0
distribution of surplus cash =($54,000) ($13,000) ($41,000) 0
Blaance 0 0 0
3rd Case The noncash assets are sold for $50,000, and any partner with a deficit is able to investup to $8,000 cash in the partnership.
Cash Assets Liabilities Dawson Feeney Hardin
Balance before liquidation $25,000 $120,000 $40,000 $31,000 $65,000 $49,000
Sale of Asset and Distribution of Loss [120000-50000 = 70000 in ratio of 30:40:30] $50,000 ($120,000**) - ($21,000) ($28,000) ($21,000)
Payment of Liability ($40,000) - ($40,000) - - -
Surplus Balance $35,000 - - $10,000 $37,000 ($12,000)
Deficit partner will bring cash up to 8000 and remaining deficit of [12000-8000 = 4000 ] will made by surplus partner in ratio of 30 :40 $8,000 - - ($1,714) ($2,286) ($8,000+$4,000**)
1714.286 2285.71429
Remaining Balance $43,000 - - $8,286 $34,714 0
Surplus distributed to Surplus Partner ($43,000) - - ($8,286) ($434,714) 0
Balance 0 0 0 0 0 0

B. Journal Entry for Case 2

Cash account ............dr $9,000

To Hardin Capital A/c............$9,000